# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

### v1

@@ B.1 Relation with the classcial SPC
[…]
@@ B.2 Proof of Proposition 1
[…]
@@ B.3 Proof of Proposition 2
$\hat{H}g_{1}\!=\!\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\!\begin{bmatrix}Q_{1}g_{1}\\ Q_{2}g_{1}\\ \mathbb{0}\\ \mathbb{0}\end{bmatrix}\!=\!H_{\textnormal{sc}}g_{1}\!=\!\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\end{bmatrix}$,
@@
$\tilde{g}_{1}$ $=H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$ (48)
$=\begin{bmatrix}Q_{1}^{{\mathsf{T}}}&Q_{2}^{{\mathsf{T}}}\end{bmatrix}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$
and $g_{1}$ can be represented as
$g_{1}=\tilde{g}_{1}+\hat{g}$
where $\hat{g}\in\textrm{null}(H_{1})=\textrm{null}(\textnormal{col}(Q_{1},Q_{2}))$ and the equality is derived from the matrix $\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}$ is invertible. Thus, the $\tilde{g}_{1}$ satisfies
[…]
$=\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}g_{1}=y_{1}.$
Since $\textnormal{col}(Q_{1},Q_{2},Q_{3},Q^{*})$ has orthonormal rows, we have
@@
and (28b) is naturally satisfied. We can further substitute $\tilde{g}_{1}$ to (28a), leading to
$\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\\ Q^{*}\end{bmatrix}\tilde{g}_{1}$ (49)
$=$ $\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ \begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}\tilde{g}_{1}\\ \end{bmatrix}=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\\ \end{bmatrix}$.
Thus, $u_{1},y_{1},\sigma_{y_{1}}$ and $\tilde{g}_{1}$ is a feasible solution for (28).
@@ B.4 Proof of Corollaries 1 and 2
As $\Gamma,\mathcal{U},\mathcal{Y}$ are convex and compact sets, $\mathcal{X}$ defined above for (22) and (28) are convex and compact. Furthermore, a quadratic function is locally Lipschitz continuous over a compact set. Thus, all conditions in Theorem 1 are satisfied and that completes the proof.
B.5 Proof of Theorem 4
The key idea for the proof of Theorem 4 is that KKT conditions of (32) and (33) have similar forms. We first illustrate that the primal and dual optimal solutions of (32) satisfy its KKT condition. Then, we present that they also satisfy the KKT condition for (33) under a specific choice of $\lambda_{g}$. That is sufficient to ensure the optimal solution of (32) is also that of (33).
We denote $x:=\textnormal{col}(\sigma_{y},u,y)$, $v:=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y)$ and we define $f(x):=\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$. Let $x^{*},g^{*}$ be the primal optimal solution of (32) and $\mu_{1}^{*},\mu_{2}^{*}$ be the dual optimal solutions for the inequality constraint and equality constraint of (32) respectively. Since (32) satisfies Slater’s condition, $x^{*},g^{*},\mu_{1}^{*}$ and $\mu_{2}^{*}$ satisfy the KKT condition and have the follow properties
$0\in\partial(f(x)+\mu_{1}^{*}(\|g\|_{1}-\alpha_{\textnormal{c}})+\mu_{2}^{*{{\mathsf{T}}}}(Hg-v^{*}))$ (50)
$\text{at}\ (x,g)=(x^{*},g^{*})\ \text{and}\ Hg^{*}=v^{*}.$
We can choose $\lambda_{g}^{*}=\mu_{1}^{*}$ and then the KKT conditions for the primal solution $\bar{x},\bar{g}$ and dual solution $\bar{\mu}$ of (33) are
$0\in\partial(f(x)+\mu_{1}^{*}\|g\|_{1}+\bar{\mu}^{{\mathsf{T}}}(Hg-v)),$ (51)
$\text{at}\ (x,g)=(\bar{x},\bar{g})\ \text{and}\ H\bar{g}=\bar{v}.$
We can let $\bar{x}=x^{*},\bar{g}=g^{*}$ and $\bar{\mu}=\mu_{2}^{*}$. Then, the KKT condition (51) holds as it becomes equivalent to (50). Thus, $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ (i.e., $(x^{*},g^{*})$) is also an optimal solution for (33) as the KKT condition is sufficient to guarantee optimality.
B.6 Proof of Proposition 5

### v2

@@ B.1 Relation with the classcial SPC
which means $u_{1},y_{1},\sigma_{y_{1}}$ is also a feasible solution of (44).
@@ B.2 Proof of Proposition 1
where we have applied the fact that $(I-\Pi_{1})g_{1}=0$ from (22b). Thus, the set of variables $u_{1},y_{1},\sigma_{y_{1}}$ and $g_{1}$ is also a feasible solution for (21).
@@ B.3 Proof of Proposition 2
$\hat{H}g_{1}\!=\!\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\!\begin{bmatrix}Q_{1}g_{1}\\ Q_{2}g_{1}\\ \mathbb{0}\\ \mathbb{0}\end{bmatrix}\!=\!H_{\textnormal{sc}}^{*}g_{1}\!=\!\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\end{bmatrix}$,
@@
$\tilde{g}_{1}$ $=H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$
$=\begin{bmatrix}Q_{1}^{{\mathsf{T}}}&Q_{2}^{{\mathsf{T}}}\end{bmatrix}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$,
so that $u_{1},y_{1},\sigma_{y_{1}}$ and $\tilde{g}_{1}$ is an optimal solution of (28). Since $\textnormal{col}(Q_{1},Q_{2},Q_{3},Q^{*})$ has orthonormal rows, we have
@@
and (28b) is naturally satisfied. Further substituting $\tilde{g}_{1}$ to (28a) leads to
$\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\\ Q^{*}\end{bmatrix}\tilde{g}_{1}$
$=$ $\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ \begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}\tilde{g}_{1}\\ \end{bmatrix}=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\\ \end{bmatrix}$,
where the second equality comes from we can represent $g_{1}$ as $g_{1}=\tilde{g}_{1}+\hat{g}$ with $\hat{g}\in\textrm{null}(H_{1})=\textrm{null}(\textnormal{col}(Q_{1},Q_{2}))$ and then
$\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}\tilde{g}_{1}=\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}g_{1}=y_{1}.\vskip-5.69054pt$
That shows the satisfaction of (28a).
@@ B.4 Proof of Corollaries 1 and 2
As $\Gamma,\mathcal{U},\mathcal{Y}$ are convex sets, $\mathcal{X}$ defined above for (22) and (28) are convex. Thus, all conditions in Theorem 1 are satisfied and that completes the proof.
B.5 Proof of Proposition 5
@@
Appendix C Additional example: Two-wheeled robot
We here present the closed-loop control performance of DDPC variants and Equality-form SPC for a two-wheeled robot to illustrate the effectiveness of the framework in a more realistic application. The kinematic model of the robot is nonlinear [li2019online]
$z_{x}(k+1)$ $=z_{x}(k)+\Delta t\cdot\cos(z_{\delta}(k))\cdot v(k),$ (50)
$z_{y}(k+1)$ $=z_{y}(k)+\Delta t\cdot\sin(z_{\delta}(k))\cdot v(k),$
$z_{\delta}(k+1)$ $=z_{\delta}(k)+\Delta t\cdot w_{t},$
where $(z_{x},z_{y},z_{\delta})$ denotes the position and heading angle of the robot (i.e., state), $v,w$ represent tangential and angular velocities (i.e., input) and $\Delta t=0.025\,\mathrm{s}$ is the sampling time. We aim to track the following heart-shaped reference trajectory $(r_{x},r_{y},r_{\delta})$ while minimizing the control effort
$r_{x}(t)=$ $16\sin^{3}(t-6),$
[…]
$r_{\delta}(t)=$ $\ \arctan\left(\frac{r_{y}(t+1)-r_{y}(t)}{r_{x}(t+1)-r_{x}(t)}\right).$
We choose the cost matrices $Q=\mathrm{diag}(2,2,8),$ and $R=6.25\times 10^{-4}$, and set the weighting parameter $\lambda_{y}=3\times 10^{6}$ for all approaches. The remaining parameters are tuned separately for each approach.
To improve the accuracy of the data-driven representation and the tracking performance, we partition the state space into different regions and construct a corresponding data-driven model for each of them. This can be viewed as a local linear approximation of the nonlinear system. Specifically, as the nonlinearity of the system mainly arises from the coupled terms $\sin(z_{\delta})\cdot v$ and $\cos(z_{\delta})\cdot v$ in states $z_{x}$ and $z_{y}$, we divide the state-space into $4$ regions according to the orientation, that are $0,\frac{\pi}{2}),[\frac{\pi}{2},\pi),[\pi,\frac{3\pi}{2})$ and $\frac{3\pi}{2},2\pi)$. We collect $4$ trajectories of length $1200$ with initial states $(0,0,\frac{\pi}{4}),(0,0,\frac{3\pi}{4}),(0,0,\frac{5\pi}{4})$ and $(0,0,\frac{7\pi}{4})$ to capture the behavior of the nonlinear system within these regions. The input signals $v$ and $w$ are sampled uniformly from $[10,20]$ and $\frac{\pi}{6},\frac{\pi}{6}]$, respectively. Then, for each approach, we construct $4$ corresponding trajectory libraries over these regions with $T_{\textnormal{ini}}=5$ and $N=12$. During the online tracking, we switch the data library based on the robot’s current orientation.
The closed-loop tracking performance of different approaches is shown in Fig. 8. All approaches are able to approximately track the prescribable reference trajectory. Among them, the L-DDPC and SPC exhibit noticeable deviation and oscillation behavior, while the C-DDPC and O-DDPC achieve superior tracking performance.
(a) L-DDPC
(b) SPC
(c) C-DDPC
(d) A-DDPC
Figure 8: Closed-loop tracking performance of the two-wheel robot with different DDPC variants.
(a) L-DDPC
(b) SPC
(c) C-DDPC
(d) A-DDPC
Figure 9: Closed-loop tracking performance of the two-wheel robot with different DDPC variants.
We further report the computation times of DDPC variants in Table 2. The optimization problems are solved using Mosek [mosek] on an Intel Core i7-9750H CPU with 16GB RAM and each controller is executed $100$ times to obtain the mean computation time. The computation times of all approaches are around $3\times 10^{-2}\,\mathrm{s}$, which are relatively small, although slightly larger than the discretization time $0.025\,\mathrm{s}$, with a gap of around $5\times 10^{-3}\,\mathrm{s}$. We can compute the control input less frequently and apply multi-step inputs from the controller between updates. Fig. 9 shows the control performance of the DDPC variants when the optimal control input is updated every $0.05\,\mathrm{s}$ and two control inputs are applied at each update. The results remain satisfactory and meet the real-time implementation requirement. Furthermore, this gap could potentially be reduced by developing a solver tailored to the specific structure of our problem.
Table 2: Computation time of DDPC variants (Unit $\mathrm{s}$)
L-DDPC C-DDPC O-DDPC SPC
Two-wheeled robot 0.030 0.032 0.029 0.0316